OpenStudy (anonymous):

how many moles in 1.2g of water? how many moles in 1.8g of MgSO4? write a probable formula for magnesium sulfate hydrate.

OpenStudy (anonymous):

medals to how ever helps

OpenStudy (frostbite):

You know the relationship between the amount of substance in moles and the mass of a chemical?

OpenStudy (anonymous):

that why i need help please

OpenStudy (frostbite):

Okay the relation can be summarized from this equation: "for a chemical \(\sf X_{a}Y_{b}\) we say that: \[\large n(\sf X_{a}Y_{b}) = \frac{ m(\sf X_{a}Y_{b}) }{ M(\sf X_{a}Y_{b}) }\] Here is: \(n\) the amount of substance in moles. \(m\) is the mass in gram \(M\) is the molar mass in gram/moles and given from the equation: \[\large M(\sf X_{a}Y_{b})=a \times M(X)+b \times M(Y)\]

OpenStudy (anonymous):

thank you